OpenStudy (anonymous):

What is the x-intercept of the line perpendicular to the line defined by 3x-2y=6 and whose y-intercept is 2?

OpenStudy (radar):

You understand that this line will have a slope that is the negative reciprocal of the defined line 3x-2y=6. Put this line equation in standard form 2y=3x-6 then y=(3/2)(x)-3 that slope is 3/2 so the negative reciprocal is - 2/3. That is the slope for the perpendicular line which you are to find the x intercept.

OpenStudy (anonymous):

I dont get it.

OpenStudy (radar):

With the defined line (3x-2y)=6 do you know what the intercepts are?

OpenStudy (anonymous):

no

OpenStudy (radar):

Those are the points on the line where they cross the axis. For the defined line you can calculate these intercepts by assiging 0 to y and then calculating x, and then assigning 0 to x and calculating y. Review the line equation and line graphing then the characteristics of perpendicular line to a given line.

OpenStudy (radar):

I have to run, maybe someone else can continue with you or repost the question.

OpenStudy (anonymous):

x=2 and y=-3?

OpenStudy (anonymous):

or is it the other way around, oh well. Thanks for helping me! :)

OpenStudy (radar):

Yes, and now that you understand the intercept thing, lets develop the equation of the perpendicular line. as shown earlier the slope is -2/3 and you know that the y intercept is 2. You have a point now known (0,2) So you can develop the equation of the perpendicular line from this: m=(y1-y2)/(x1-x2) or: -2/3= (2-y)/(0-x) putting into slope intercept form you will obtain: y=(-2/3)x+2 Now you got your intercepts.......When x=0,Y=2 When y=0,x=3

OpenStudy (radar):

x intercept is 3